# Properties of neutron stars and strangeness-mixed stars from a pion mean-field approach

### v1

@@ Abstract
We investigate the properties of the static neutron stars and strangeness-mixed stars, based on the equations of state derived from a pion mean-field approach. Using the empirical data on the pion-nucleus scattering and bulk properties of nuclear matter, we have already fixed all the parameters in a previous work, where the nucleons and hyperons were shown to be modified in various nuclear medium. In the current work, we first examine the energy and pressure inside a neutron star. We show that the central densities in various neutron stars vary within the range of $(3-6)\rho_{0}$, where $\rho_{0}$ is the normal nuclear matter density. The mass-radius relations are obtained and discussed. As the slope parameter for neutron matter increases, the radii of the neutron stars increase with their masses fixed. We also study the strangeness-mixed stars or the hyperon stars using the same sets of the parameters. As the strangeness content of strange matter increases, the binding energy per nucleon is saturated and the corresponding equation of state becomes softened. Consequently, the central densities of the strangeness-mixed stars increase. Assuming that recently observed neutron stars are the strangeness-mixed ones, we find that the central densities increase. In the case of the pure strange stars, the central densities reach almost $(5-6)\rho_{0}$.
@@ I Introduction
The current work is organized as follows: In Sec. II, we recapitulate the pion mean-field approach, focusing on the collective Hamiltonian and a medium modification of the model using the variational approach. In Sec. III, we examine the mass, energy, and pressure densities, which will be used in the Tolman-Oppenheimer-Volkoff (TOV) equation. In Sec. IV, we discuss the results for the masses and radii of the neutron stars, and then we extend the formalism to investigate the hybrid stars, employing the strange matter. In Sec. V, we draw conclusions and summarize the current work.
@@ II Pion mean-field approach
The dynamical parameters will be determined by using the experimental and empirical data. This method is often called “a model-independent” approach for the chiral soliton, which has a virtue when it is extended to the investigation of bulk properties of nuclear matter and baryon properties in it.
Since the classical soliton does not carry quantum numbers of baryons, we need to quantize it. As mentioned already, the quantum fluctuations of the pion fields are suppressed by the large $N_{c}$ approximation. On the other hand, the fluctuation of the pion field to the zero-mode directions, which are always related to translational and rotational symmetries, must be considered. Rotational zero modes determine the quantum numbers of baryons. For details, we refer to Refs. [31, 32]. Having performed the zero-mode quantization, we derive the collective Hamiltonian [34, 36, 35]
@@
where the $I_{1,2}$ are the moments of inertia and $\hat{J}_{a}$ ($a=1,\cdots 8$) are the generators of the the $\mathrm{SU(3)}$ group, which are directly related to the right angular velocities $\Omega_{a}$. The eighth component of $J_{a}$ is constrained to be
@@
where $B$ is the baryon number. $J_{8}$ is related to the right hypercharge $Y_{R}=-Y^{\prime}=(2/\sqrt{3})J_{8}$. While in the Skyrme model $J_{8}$ is constrained by the Wess-Zumino term, the presence of the $N_{c}$ valence quarks constrained it within the current approach. The constraint (10) has significant physical implications: It selects the lowest allowed representations. If the baryon multiplet with $Y=1$ consists of $12J+$ 1 multiplet, the spin of the multiplet is set to be $J$. Thus, the baryon octet and decuplet are selected. In the representation $\mathcal{R}=(p,q)$, the eigenvalues of the Hamiltonian $H_{\mathrm{rot}}$ in Eq. (9) are given as
@@
where the $\alpha$, $\beta$, $\gamma$ expressed in terms of the moments of inertia
@@
The dynamical parameters $\alpha$, $\beta$, and $\gamma$ can be evaluated within the self-consistent SU(3) $\chi$QSM. However, we will adopt the model-independent approach, as previously mentioned. So, we will fix all dynamical parameters by using the experimental data rather than by computing within the model [77, 36]. Furthermore, we employ the variational method to modify the collective Hamitonian (8) in nuclear medium. We briefly recapitulate the general formalism for the medium modification of $H$. For details, we refer to Refs. [72, 73].
@@
×$\times\frac{5(M_{cl}-M_{\mathrm{cl}}^{\ast}+E_{(1,1)1/2}-E_{(1,1)1/2}^{*})}{3(m_{s}-\hat{m})},$ (29)
@@
where $a_{\mathrm{sym}}$ denotes the nuclear symmetry energy at saturation point, $L_{\mathrm{sym}}$ stands for that of its slope parameter $L_{\mathrm{sym}}$, and $K_{\mathrm{sym}}$ represents the asymmetric part of the incompressibility. They are expressed as
@@ III Energy densities and pressures for neutron stars
To study the mass-radius relation of the neutron stars, we consider static and spherically symmetric non-rotating neutron stars, which satisfies the Tolman-Oppenheimer-Volkoff (TOV) equation. The mass function of the neutron star $M(r)$ is given as a function of radius $r$
@@ IV.1 Neutron stars
Figure 3: Masses of the neutron star are drawn as functions of the dimensionless central density $\lambda=\rho/\rho_{0}$ with $\rho_{0}=0.16\,\mathrm{fm}^{-3}$ in unit of the solar mass $M_{\odot}$. The red bar corresponds to the data from Ref. [87]. The data for the orange, blue, and green bars are taken from Refs. [14, 12, 13].
In Fig. 3, we show the masses of the neutron star in units of the solar mass as functions of dimensionless central density $\lambda$ with the four different parameter sets used. While the results are not sensitive to $a_{\mathrm{sym}}$, they depend on $L_{\mathrm{sym}}$ in lower nuclear matter density. The results indicate that the central density of the neutron stars observed in Refs. [14, 12, 13] is approximately $\lambda\approx 4$ or $\rho\approx 0.64\,\mathrm{fm}^{-3}$. On the other hand, The central density for the neutron star from GW190814 [87] is around $\lambda\approx 6$ or $\rho\approx 0.96\,\mathrm{fm}^{-3}$. The masses of the neutron stars are almost saturated at around $M\approx 2.5M_{\odot}$.
@@
Figure 4: Mass-radius relations of neutron stars. The black curves represent the results with the four different sets of parameters given in Table 1. The value of the incompressibility is chosen as $K_{0}=240$ MeV. The red and green lines with stars are taken from the LIGO experimental data [90]. The red bar corresponds to the data from Ref. [87]. The data for the orange, blue, and green bars are taken from Ref. [14, 12, 13].
Figure 4 depicts the mass-radius relations for the neutron stars. As already discussed above, the mass-radius relations are not sensitive to $a_{\mathrm{sym}}$, the radius tends to increase as the slope parameter $L_{\mathrm{sym}}$ increases, in particular, when the neutron matter density is low. As the density increases, the dependence of the mass-radius relations on $L_{\mathrm{sym}}$ becomes smaller. Figure 4 also indicates that the central densities in PSR J0740+6620, PSR J0348+0432, and PSR J1614+2330 [14, 12, 13] are estimated to be around $3.5\rho_{0}$. The current calculation estimates the central density of the neutron star from GW190814 [87] to be around $6.5\rho_{0}$. When $L_{\mathrm{sym}}=50$ MeV is used, the radii of the neutron stars are approximately 12 km except for that from the neutron star corresponding to GW190814, of which the radius is around 11 km. On the other hand, if we employ $L_{\mathrm{sym}}=60$ MeV, the radii of the neutron stars tend to increase, i.e. $R\approx 12.5$ km, again except for that from GW190814, which remains almost stable as $L_{\mathrm{sym}}$ is changed.
Figure 5: Dependence of the neutron star mass-radius relation on the incompressibility, $K_{0}$. The values of the symmetry energy at saturation density and the corresponding slop parameter are chosen as $a_{\rm sym}=32$ MeV and $L_{\rm sym}=60$ MeV, respectively. The observational and empirical data are the same as in Fig. 4.
The empirical values of the nuclear matter incompressibility $K_{0}$ are given in a wide range. For example, the value of $K_{0}$ was often predicted as $K_{0}\approx(290\pm 70)$ MeV [91, 92, 93, 94, 95, 96]. In Ref. [97], however, a smaller value was extracted from the data on the isoscalar giant monopole resonance in ${}^{208}\mathrm{Pb}$, based on relativistic mean-field models: $K_{0}=(248\pm 8)$ MeV. In Ref. [98], $K_{0}=(240\pm 20)$ MeV was extracted from a reanalysis of the data on the energies of the giant monopole resonances in even-even 112-124Sn and 106,100-116Cd and earlier data on $58\leq A\leq 208$ nuclei. A similar value was also estimated in Ref. [99]. Thus, while we choose $K_{0}=240\,\mathrm{MeV}$, it is of great interest to examine the dependence of the mass-radius relations of the neutron stars on $K_{0}$. In Fig. 5, we demonstrate the mass-radius relations of the neutron stars, varying the incompressibility of nuclear matter $K_{0}$ from 210 MeV to 260 MeV. We take the values of $a_{\mathrm{sym}}=32$ MeV and $L_{\mathrm{sym}}=60$ MeV. The tendency of the mas-radius relations is obvious in general. If we increase $K_{0}$ from 220 MeV to 260 MeV, the mass-radius relations remain as almost the same shape but are shifted to the right. So, the radii of the neutron stars get larger as $K_{0}$ increases.
@@ IV.2 Strageness-mixed stars
In Fig. 6, we draw the results for the energy and pressure as functions of $\lambda$ with the sum of the parameters $\delta_{s}$ changed from 0 to 1. While $\sum\delta_{s}=0$ corresponds to pure neutron matter, $\sum\delta_{s}=1$ indicates pure strange matter. The left panel of Fig. 6 shows that as the number of the strange quarks increases, the energy per nucleon starts to decrease. When $\lambda$ reaches about $0.16$, $\varepsilon(\lambda)$ becomes negative and turns positive, which exhibits a behavior similar to normal nuclear matter. As the content of the strangeness increases, the saturation point gets lower till it reaches $-5.97$ MeV at $\lambda\simeq 1$ for pure strange matter. However, strangeness-mixed matter has a shallow saturation point in comparison with nuclear matter [72]. Consequently, the pressure of strangeness-mixed matter becomes weaker as the content of the strangeness increases, as demonstrated in the right panel of Fig. 6. Interestingly, the pressure is lessened by about four times for pure strange matter. This implies that the EoS for strangeness-mixed matter is softer than neutron matter, as discussed in Refs. [100, 101, 102].
@@ V Summary and outlook
We have investigated the properties of the neutron stars and strangeness-mixed stars, using the equation of the state for both neutron matter and strangeness-mixed matter, which were derived from the medium-modified pion mean-field approach (chiral quark-soliton model) [72]. All the relevant parameters were already fixed in Ref. [72]. A great virtue of this approach is that the medium modifications of the nucleon and hyperons have been considered, in addition to describing the bulk properties of nuclear media. The results for the mass-radius relations were well described in the current work. The central densities in PSR J0740+6620, PSR J0348+0432, and PSR J1614+2330 [14, 12, 13] were obtained to be around $3.5\rho_{0}$, while the central density corresponding to the neutron star from GW190814 [87] was estimated to be $6.5\rho_{0}$. We also examined the dependence of the mas-radius relations on the incompressibility $K_{0}$. As $K_{0}$ increases, central density becomes lower and the radius increases. To investigate the strangeness-mixed stars, we first examined the equation of state for strangeness-mixed matter, as the strangeness content varied from 0 to 1, which correspond to neutron matter and pure strange matter, respectively. At $\lambda\approx 0.16$, the binding energy becomes negative and turns positive, which behaves as normal nuclear matter. As the content of the strangeness increases, the binding energy becomes further lower, though strange matter has a shallow saturation point in comparison with nuclear matter. As a result, the pressure of strange-mixed matter becomes lessened as the strangeness content increases. Note that the pressure is increased by about four times for pure strange matter. This implies that the EoS for strangeness-mixed matter is softened. Thus, the masses of the strangeness-mixed star become smaller at the same central density. We conclude that the strange stars are more compact than the neutron stars.

### v2

@@ Abstract
We investigate the properties of the static neutron stars and strangeness-mixed stars, based on the equations of state derived from a pion mean-field approach. Using the empirical data on the pion-nucleus scattering and bulk properties of nuclear matter, we have already fixed all the parameters in a previous work, where the nucleons and hyperons were shown to be modified in various nuclear medium. In the current work, we first examine the energy and pressure inside a neutron star. We show that the central densities in various neutron stars vary within the range of $(3-4)\rho_{0}$, where $\rho_{0}$ is the normal nuclear matter density. The mass-radius relations are obtained and discussed. As the slope parameter for neutron matter increases, the radii of the neutron stars increase with their masses fixed. We also study the strangeness-mixed stars or the hyperon stars using the same sets of the parameters. As the strangeness content of strange matter increases, the binding energy per nucleon is saturated and the corresponding equation of state becomes softened. Consequently, the central densities of the strangeness-mixed stars increase. Assuming that recently observed neutron stars are the strangeness-mixed ones, we find that the central densities increase.
@@ I Introduction
Since we assume that neutron stars have already been cooled and stationary, we use the Tolman-Oppenheimer-Volkoff (TOV) equation for stationary neutron stars. In addition, we need to consider the effects due to the presence of leptons [74, 75, 76, 77, 78, 79, 80, 81], which brings about the contribution from the protons, in particular, at the crust of neutron stars. This will come into play to restrict the EoS [82].
The current work is organized as follows: In Sec. II, we recapitulate the pion mean-field approach, focusing on the collective Hamiltonian and a medium modification of the model using the variational approach. In Sec. III, we examine the mass, energy, and pressure densities, which will be used in the TOV equation. We will also show how we can include leptonic contributions. In Sec. IV, we discuss the results for the masses and radii of the neutron stars, and then we extend the formalism to investigate the hybrid stars, employing the strange matter. In Sec. V, we draw conclusions and summarize the current work.
@@ II Pion mean-field approach
The dynamical parameters will be determined by using the experimental and empirical data. This method is often called “a model-independent” approach for the chiral soliton, which has a virtue when it is extended to the investigation of bulk properties of nuclear matter and baryon properties in it. Since the classical soliton does not carry quantum numbers of baryons, we need to quantize it. As mentioned already, the quantum fluctuations of the pion fields are suppressed by the large $N_{c}$ approximation. On the other hand, the fluctuation of the pion field to the zero-mode directions, which are always related to translational and rotational symmetries, must be completely treated. Rotational zero modes determine the quantum numbers of baryons such as the spin and isospin. For details, we refer to Refs. [31, 32].
Having performed the rotational zero-mode quantization, we derive the collective Hamiltonian [34, 36, 35]
@@
where the $I_{1,2}$ are the moments of inertia and $\hat{J}_{a}$ ($a=1,\cdots 8$) are the generators of the $\mathrm{SU(3)}$ group, which are directly related to the right angular velocities $\Omega_{a}$. The eighth component of $J_{a}$ is constrained to be
@@
[…]
@@
where $\alpha$, $\beta$, $\gamma$ are expressed in terms of the moments of inertia
@@
The dynamical parameters $\alpha$, $\beta$, and $\gamma$ can be evaluated within the self-consistent SU(3) $\chi$QSM. However, we will adopt the model-independent approach, as previously mentioned. So, we will fix all dynamical parameters by using the experimental data rather than by computing within the model [86, 36]. Furthermore, we employ the variational method to modify the collective Hamiltonian (8) in nuclear medium. We briefly recapitulate the general formalism for the medium modification of $H$. For details, we refer to Refs. [72, 73].
Before we proceed to discuss the density dependence of the binding energy per nucleon, we will briefly mention the corrections from the translational zero modes that make the nucleon carry the linear momentum. The zero-mode translation can be included by translating the chiral soliton as
$U_{x}(\bm{x}-\bm{X}(\tau))=T_{\bm{X}(\tau)}U_{c}(\bm{x})T_{\bm{X}(\tau)}^{\dagger},$ (15)
where $\bm{X}(\tau)$ is the translational zero modes conjugate to the linear momentum of the soliton. $T_{\bm{X}(\tau)}$ is a unitary SU(2) matrix that governs the corresponding translation. Once we carry out the quantization by expanding the effective chiral action with respect to the velocity $\dot{\bm{X}(\tau)}$ to the second order, and compute the functional integration over translational zero modes, we have kinetic corrections to the collective Hamiltonian given in Eq. (8)
$H^{\prime}=H+\frac{\hat{\bm{P}}^{2}}{2M_{\mathrm{cl}}},$ (16)
where the kinetic term is of the order $\mathcal{O}(N_{c}^{0})$ in the large $N_{c}$ limit. $\hat{\bm{P}}$ is the linear momentum operator conjugate to $\hat{\bm{X}}$. However, there is one important caveat. The classical mass $M_{\mathrm{cl}}$ already contains spurious contributions from the center-of-mass motion, which must be subtracted. This process is nontrivial as was discussed in depth within the Skyrme models [87, 88], with $1/N_{c}$ pion-loop corrections included. Technically, this has resemblance with the random phase approximation in many-body theory. We may expect that two main subtraction terms expressed as
$\Delta E=-\frac{\langle P^{2}\rangle}{2M_{\mathrm{cl}}}-\frac{\langle J^{2}\rangle}{2I},$ (17)
where $\langle p^{2}\rangle$ and $\langle J^{2}\rangle$ denote the average values of $P^{2}$ and $J^{2}$. The derivation of these values is beyond the current theoretical framework. Thus, we assume that the classical mass of the nucleon effectively carry the corrections from the kinetic energy.
@@
×$\hskip-28.45274pt\times\frac{5(M_{cl}-M_{\mathrm{cl}}^{\ast}+E_{(1,1)1/2}-E_{(1,1)1/2}^{*})}{3(m_{s}-\hat{m})},$ (32)
@@
where $a_{\mathrm{sym}}$ denotes the nuclear symmetry energy at the saturation point, $L_{\mathrm{sym}}$ stands for that of its slope parameter $L_{\mathrm{sym}}$, and $K_{\mathrm{sym}}$ represents the asymmetric part of the incompressibility. They are expressed as
@@ III Energy densities and pressures for neutron stars
To study the mass-radius relation of the neutron stars, we consider static and spherically symmetric non-rotating neutron stars, which satisfy the TOV equation. The mass function of the neutron star $M(r)$ is given as a function of radius $r$
@@
Although the leptons inside a neutron star do not participate in the strong interaction, this will cause certain contributions from the protons in the crust of the neutron star. We use a relativistic Fermi gas model to describe the electrons inside a the neutron star [98, 76], so that we obtain the energy density and pressure of the electron gas as
$\mathcal{\epsilon}_{e}$ $=\frac{3}{4}n{e}\mu_{e},\;\;\;P_{\mathrm{e}}=\frac{1}{4}n_{e}\mu_{e},$ (48)
where $n_{e}$ denotes the number density of the electrons, and $\mu_{e}$ is the chemical potential of the electron.
Since we consider the beta-stable condition, the chemical potential of the electron can be expressed in terms of the symmetry energy as [78]
$\mu_{e}=4\delta E_{\mathrm{sym}}(\lambda).$ (49)
Since the neutron star satisfies the charge neutrality condition, the electron number density must be the same as the proton number density, which brings about the contribution from the protons. Thus, the symmetry energy depends not only on the density of the neutron star but also on the fraction of the protons. We fix the electron and proton fractions to be $n_{e}=n_{p}=0.1$ in the current work, assuming that the leptonic effects mainly come from the crust of the neutron star. So, the electron affects only the subnuclear density region, i.e., $\lambda<1$ [78].
@@ IV Results and discussions
Before we proceed to discuss the results, we want to emphasize that the current theoretical framework is based on linear density approximation. Moreover, we do not include quark matter in the current work. It implies that we can only describe physical observables in medium approximately up to $\rho\approx 4\rho_{0}$. If the medium density becomes larger than $4\rho_{0}$, the speed of sound would break the causality. Thus, in the forthcoming discussions, we will restrict ourselves to lower densities than $4\rho_{0}$.
@@ IV.1 Neutron stars
Figure 3: Masses of the neutron star are drawn as functions of the dimensionless central density $\lambda=\rho/\rho_{0}$ with $\rho_{0}=0.16\,\mathrm{fm}^{-3}$ in unit of the solar mass $M_{\odot}$. The data for the orange, blue, and green bars are taken from Refs. [14, 12, 13].
In Fig. 3, we show the masses of the neutron star in units of the solar mass as functions of dimensionless central density $\lambda$ with the four different parameter sets used. While the results are not sensitive to $a_{\mathrm{sym}}$, they depend on $L_{\mathrm{sym}}$ in lower nuclear matter density. The results indicate that the central density of the neutron stars observed in Refs. [14, 12, 13] is approximately $\lambda\approx 4$ or $\rho\approx 0.64\,\mathrm{fm}^{-3}$. However, when heavy neutron stars are considered, quark matter must come into a significant role, which we did not include in the current work. This, we present the results for the masses of neutron stars up to $\lambda\approx 4$.
@@
Figure 4: Mass-radius relations of neutron stars. The black curves represent the results with the four different sets of parameters given in Table 1. The value of the incompressibility is chosen as $K_{0}=240$ MeV. The red and green lines with stars are taken from the LIGO experimental data [101]. The data for the orange, blue, and green bars are taken from Ref. [14, 12, 13].
Figure 4 depicts the mass-radius relations for the neutron stars. As already discussed above, the mass-radius relations are not sensitive to $a_{\mathrm{sym}}$, the radius tends to increase as the slope parameter $L_{\mathrm{sym}}$ increases, in particular, when the neutron matter density is low. As the density increases, the dependence of the mass-radius relations on $L_{\mathrm{sym}}$ becomes smaller. Figure 4 also indicates that the central densities in PSR J0740+6620, PSR J0348+0432, and PSR J1614+2330 [14, 12, 13] are estimated to be around $3.5\rho_{0}$. As mentioned previously, by including the leptonic effects, the radius of the neutron stars tend to decrease. They mainly contribute to lower-density regions.
Figure 5: Dependence of the neutron star mass-radius relation on the incompressibility, $K_{0}$. The values of the symmetry energy at saturation density and the corresponding slope parameter are chosen as $a_{\rm sym}=32$ MeV and $L_{\rm sym}=60$ MeV, respectively. The observational and empirical data are the same as in Fig. 4.
The empirical values of the nuclear matter incompressibility $K_{0}$ are given in a wide range. For example, the value of $K_{0}$ was often predicted as $K_{0}\approx(290\pm 70)$ MeV [102, 103, 104, 105, 106, 107]. In Ref. [108], however, a smaller value was extracted from the data on the isoscalar giant monopole resonance in ${}^{208}\mathrm{Pb}$, based on relativistic mean-field models: $K_{0}=(248\pm 8)$ MeV. In Ref. [109], $K_{0}=(240\pm 20)$ MeV was extracted from a reanalysis of the data on the energies of the giant monopole resonances in even-even 112-124Sn and 106,100-116Cd and earlier data on $58\leq A\leq 208$ nuclei. A similar value was also estimated in Ref. [110]. Thus, while we choose $K_{0}=240\,\mathrm{MeV}$, it is of great interest to examine the dependence of the mass-radius relations of the neutron stars on $K_{0}$. In Fig. 5, we demonstrate the mass-radius relations of the neutron stars, varying the incompressibility of nuclear matter $K_{0}$ from 210 MeV to 260 MeV. We take the values of $a_{\mathrm{sym}}=32$ MeV and $L_{\mathrm{sym}}=60$ MeV. The tendency of the mass-radius relations is obvious in general. If we increase $K_{0}$ from 220 MeV to 260 MeV, the mass-radius relations remain as almost the same shape but are shifted to the right. So, the radii of the neutron stars get larger as $K_{0}$ increases.
@@ IV.2 Strageness-mixed stars
In Fig. 6, we draw the results for the energy and pressure as functions of $\lambda$ with the sum of the parameters $\delta_{s}$ changed from 0 to 1. While $\sum\delta_{s}=0$ corresponds to pure neutron matter, $\sum\delta_{s}=1$ indicates pure strange matter. The left panel of Fig. 6 shows that as the number of the strange quarks increases, the energy per nucleon starts to decrease. When $\lambda$ reaches about 0.160.160.16, $\varepsilon(\lambda)$ becomes negative and turns positive, which exhibits a behavior similar to normal nuclear matter. As the content of the strangeness increases, the saturation point gets lower till it reaches 5.97-5.97- 5.97 MeV at $\lambda\simeq 1$ for pure strange matter. However, strangeness-mixed matter has a shallow saturation point in comparison with nuclear matter [72]. Consequently, the pressure of strangeness-mixed matter becomes weaker as the content of the strangeness increases, as demonstrated in the right panel of Fig. 6. Interestingly, the pressure is lessened by about four times for pure strange matter. This implies that the EoS for strangeness-mixed matter is softer than neutron matter, as discussed in Refs. [111, 112, 113].
@@ V Summary and outlook
We have investigated the properties of the neutron stars and strangeness-mixed stars, using the equation of the state for both neutron matter and strangeness-mixed matter, which were derived from the medium-modified pion mean-field approach (chiral quark-soliton model) [72]. All the relevant parameters were already fixed in Ref. [72]. A great virtue of this approach is that the medium modifications of the nucleon and hyperons have been considered, in addition to describing the bulk properties of nuclear media. In addition, we have considered the leptonic effects, which mainly come from the crust of the neutrons stars. The results for the mass-radius relations were well described in the current work. The central densities in PSR J0740+6620, PSR J0348+0432, and PSR J1614+2330 [14, 12, 13] were obtained to be around $3.5\rho_{0}$.
We also examined the dependence of the mass-radius relations on the incompressibility $K_{0}$. As $K_{0}$ increases, central density becomes lower and the radius increases. To investigate the strangeness-mixed stars, we first examined the equation of state for strangeness-mixed matter, as the strangeness content varied from 0 to 1, which correspond to neutron matter and pure strange matter, respectively. At $\lambda\approx 0.16$, the binding energy becomes negative and turns positive, which behaves as normal nuclear matter. As the content of the strangeness increases, the binding energy becomes further lower, though strange matter has a shallow saturation point in comparison with nuclear matter. As a result, the pressure of strange-mixed matter becomes lessened as the strangeness content increases. Note that the pressure is increased by about four times for pure strange matter. This implies that the EoS for strangeness-mixed matter is softened. Thus, the masses of the strangeness-mixed star become smaller at the same central density. We conclude that the strange stars are more compact than the neutron stars.